A hot-air balloon is descending at a rate of 2.0 m/s when a passenger drops a camera.

a)If the camera is 44 m above the ground when it is dropped, how much time does it take for the camera to reach the ground?

b)If the camera is 44 m above the ground when it is dropped, what is its velocity just before it lands? Let upward be the positive direction for this problem.
Answer #1

from the kinematic equation

s= ut + ( 1/ 2) a t ^ 2

                         44= 2 t+ 4.9 t ^ 2

4.9 t ^ 2+ 2t - 44 = 0

solving quadratic equation

t = 2.79 s

Let the velocity just before collision be vthen

v = u + at

    = 2.0 m/s + 9.8 m/s^2 ( 2.79 s)

   =29.342 m/s (Answer)

I hope help you !!